# Hierarchical Clustering for Conditional Diffusion in Image Generation

### v1

@@ C.1 Generations on MNIST
Figure 8 presents an additional plot similar to those in Figure 4 and Figure 5 from the main text. This plot illustrates the generated images of the TreeDiffusion model when trained on the MNIST dataset. In each of these plots, we display randomly generated images for each cluster. Below each set of leaf-specific images, we provide a normalized histogram showing the distribution of predicted classes by an independent ResNet-50 classifier that has been pre-trained on the training data of the respective dataset. This visualization helps in understanding how well the model can generate distinct and meaningful clusters in the context of different datasets.
@@
Figure 8: TreeDiffusion model trained on MNIST. For each cluster, random newly generated images are displayed. Below each set of images, a normalized histogram (ranging from 0 to 1) shows the distribution of predicted classes from an independent, pre-trained classifier on MNIST for all newly generated images in each leaf with a significant probability of reaching that leaf.
C.2 Additional Generation Examples for TreeVAE vs. TreeDiffusion
@@
Figure 9: For each example, we show image generations from every leaf of the (top) TreeVAE and (bottom) TreeDiffusion model, both trained on the MNIST dataset. Each row shows the generated images from all leaves of the respective model, starting with the same root sample. The corresponding leaf probabilities are shown above each image and are identical across both models by design.
@@
Figure 10: For each example, we show image generations from every leaf of the (top) TreeVAE and (bottom) TreeDiffusion model, both trained on the FashionMNIST dataset. Each row shows the generated images from all leaves of the respective model, starting with the same root sample. The corresponding leaf probabilities are shown above each image and are identical across both models by design.
@@
Figure 11: For each example, we show image generations from every leaf of the (top) TreeVAE and (bottom) TreeDiffusion model, both trained on the CIFAR-10 dataset. Each row shows the generated images from all leaves of the respective model, starting with the same root sample. The corresponding leaf probabilities are shown above each image and are identical across both models by design.
@@
Figure 12: For each example, we show image generations from every leaf of the (top) TreeVAE and (bottom) TreeDiffusion model, both trained on the CUBICC dataset. Each row shows the generated images from all leaves of the respective model, starting with the same root sample. The corresponding leaf probabilities are shown above each image and are identical across both models by design.
Appendix D Implementation Details
@@
# of latent channels 16 64 64
# of bottom-up channels 32 128 128
@@
# of initial epochs 150 150 150
# of smalltree epochs 150 150 150
# of intermediate epochs 80 0 0
# of fine-tuning epochs 200 0 0
@@
We utilize a modified variant of the TreeVAE model, employing CNNs for its operations. The representations within TreeVAE are maintained in a 3-dimensional format, where the first dimension signifies the number of channels, while the subsequent two dimensions denote the spatial dimensions of the representations. Consequently, the deterministic variables in the bottom-up pathway possess a dimensionality of (# of bottom-up channels, representation dimension, representation dimension), while the stochastic variables in the top-down tree structure have a dimensionality of (# of latent channels, representation dimension, representation dimension). Table 4 provides details on the remaining parameters utilized for the TreeVAE model and its training. For more information, please refer to the code.

### v2

@@ 0.C.1 Generations on CIFAR-10
Figure˜8 presents an additional plot similar to the one for FashionMNIST from the main text. This plot illustrates the generated images of the TreeDiffusion model when trained on the CIFAR-10 dataset. In each of these plots, we display randomly generated images for each cluster. Below each set of leaf-specific images, we provide a normalized histogram showing the distribution of predicted classes by an independent ResNet-50 classifier that has been pre-trained on the training data of the dataset. This visualization helps to understand how well the model can generate distinct and meaningful clusters in the context of different datasets.
@@
Figure 8: TreeDiffusion model trained on CIFAR-10. For each cluster, random newly generated images are displayed. Below each set of images, a normalized histogram (ranging from 0 to 1) shows the distribution of predicted classes from an independent, pre-trained classifier on MNIST for all newly generated images in each leaf with a significant probability of reaching that leaf.
0.C.2 Additional Generation Examples for TreeVAE vs. TreeDiffusion
@@
Figure 9: For each example, we show image generations from every leaf of TreeVAE, TreeVAE + Diffusion, and TreeDiffusion model, all trained on MNIST. Each row shows the generated images from all leaves of the respective model, starting with the same root sample.
@@
Figure 10: For each example, we show image generations from every leaf of TreeVAE, TreeVAE + Diffusion, and TreeDiffusion model, all trained on FashionMNIST. Each row shows the generated images from all leaves of the respective model, starting with the same root sample.
@@
Figure 11: For each example, we show image generations from every leaf of TreeVAE, TreeVAE + Diffusion, and TreeDiffusion model, all trained on CIFAR-10. Each row shows the generated images from all leaves of the respective model, starting with the same root sample.
@@
Figure 12: For each example, we show image generations from every leaf of TreeVAE, TreeVAE + Diffusion, and TreeDiffusion model, all trained on CUBICC. Each row shows the generated images from all leaves of the respective model, starting with the same root sample.
Appendix 0.D Implementation Details
@@
Latent channels 16 64 64
Bottom-up channels 32 128 128
@@
Initial epochs 150 150 150
Smalltree epochs 150 150 150
Intermediate epochs 80 0 0
Fine-tuning epochs 200 0 0
@@
We use a modified variant of the TreeVAE model, employing CNNs for its operations. The representations within TreeVAE are maintained in a 3D format, where the first dimension signifies the number of channels, while the subsequent two dimensions denote the spatial dimensions of the representations. Consequently, the deterministic variables in the bottom-up pathway possess a dimensionality of (Nb. of bottom-up channels, representation dimension, representation dimension), while the stochastic variables in the top-down tree structure have a dimensionality of (Nb. of latent channels, representation dimension, representation dimension). Table 4 provides details on the remaining parameters utilized for the TreeVAE model and its training. For more information, please refer to the code.